OpenStudy (anonymous):

Anyone good at calculus 3?

OpenStudy (anonymous):

Here is the question: Find the maximum value of the function f(x, y, z) = x + 5y + 11z on the curve of intersection of the plane x - y + z = 1 and the cylinder x^2 + y^2 = 1.

OpenStudy (anonymous):

differential equations ?

OpenStudy (anonymous):

No calc 3

OpenStudy (anonymous):

hmmm

OpenStudy (kainui):

Do you know what lagrange multipliers are?

OpenStudy (anonymous):

I have no idea enlighten me please